# Toward Stable Semi-Supervised Remote Sensing Segmentation via Co-Guidance and Co-Fusion

### v1

@@ IV-B Implementation Details
IV-C Comparison with State-of-the-Art Methods
We report comparison results across six widely used remote-sensing benchmarks to evaluate the performance of the proposed Co2S comprehensively. We conduct extensive comparisons against the supervised baseline (“OnlySup”), which is trained exclusively on the labeled data, as well as several general semi-supervised baselines, including FixMatch [sohn2020fixmatch], UniMatch [yang2023revisiting], U2PL [wang2022semi], and remote-sensing-specific approaches like WSCL [lu2023weak], DWL [huang2024decouple], and MUCA [wang2025semi].
IV-C1 Quantitative Comparison and Analysis
The comprehensive comparisons across six diverse benchmarks (Tables I-VI) demonstrate that Co2S consistently achieves leading performance, particularly in regimes with extremely scarce annotations. On the WHDLD [shao2020multilabel] dataset, Co2S outperforms the strong competitor UniMatch by 3.7% under the extremely sparse 1/24 split, demonstrating its capability to parse diverse land-cover features accurately. Similarly, on the LoveDA [wang2021loveda] dataset, the method yields a remarkable 12.3% improvement over the supervised baseline at the 1/40 ratio, showing superior performance against significant urban-rural distribution shifts. For the ultra-high-resolution Potsdam [isprs-annals-I-3-293-2012] dataset, Co2S achieves the best performance at the 1/32 split with 74.3% mIoU, surpassing remote-sensing-specific designs like DWL. Extending to the large-scale GID-15 [tong2020land] dataset, our method outperforms UniMatch by 1.5% at the 1/8 split, indicating proficiency in handling diverse land-cover characteristics across vast spatial areas. Furthermore, in the extraterrestrial environments of MER and MSL [li2022stepwise], Co2S delivers the highest accuracy on these benchmarks characterized by extreme class imbalance. Notably, on MSL, it boosts the mIoU by 2.3% over the runner-up DWL at the 1/4 ratio, showing significant gains in identifying challenging targets such as “tracks” where it exceeds the second-best UniMatch by 5.2%. Crucially, a consistent trend observed across all benchmarks is that the performance advantage of Co2S over baselines amplifies as the labeled ratio decreases. This phenomenon validates the stability of our Co2S, confirming that the synergistic foundation model priors effectively prevent model collapse and sustain high accuracy even in the most extreme low-data annotation regimes.
IV-C2 Qualitative Visualization Analysis
To further substantiate these quantitative findings, we provide a visual comparison of segmentation results across all six datasets in Fig. 4. As observed, baseline methods frequently struggle with semantic confusion in complex scenes. For instance, in the Potsdam dataset (Row f), UniMatch [yang2023revisiting] and DWL [huang2024decouple] incorrectly classify large areas of impervious surface (white) as clutter (red). Similarly, in WHDLD (Row a), the vegetation (green) in the top-left corner is misidentified as water (blue) by most competitors. Moreover, our method demonstrates superior capability in handling fine-grained objects and texture details. In the MER dataset (Row c), baselines suffer from severe over-segmentation, generating excessive false positives for scattered small rocks (purple), whereas Co2S precisely suppresses this noise to match the Ground Truth. Likewise, in MSL (Row d), other methods tend to misclassify large areas of bedrock (blue) as gravel (olive green), while Co2S accurately preserves the correct material distribution. These visualizations confirm that our framework effectively harmonizes global semantic consistency with local structural precision across diverse and challenging remote sensing environments, successfully rectifying category confusion while preserving fine-grained details.
@@ IV-C3 Pseudo-label Quality and Stability
where $\hat{\mathbf{y}}_{u}$ is the pseudo-label, $\mathbf{y}_{gt}$ is the ground truth, and $c$ is the confidence score. The evolution curves for the first 10 epochs on the WHDLD dataset (1/24 split) are plotted in Fig. 5. The results reveal distinct training dynamics: 1) The supervised baseline (OnlySup) exhibits extreme volatility, oscillating drastically between 50% and 90%. This highlights the instability of learning from scratch with scarce annotations. 2) Standard methods like FixMatch [sohn2020fixmatch] and UniMatch [sohn2020fixmatch] show signs of early saturation. They quickly plateau at a sub-optimal accuracy level ($\sim$80%-88%), suggesting that they are trapped by confirmation bias early in the training process and fail to correct initial errors. 3) In contrast, Co2S (red curve) demonstrates immediate semantic alignment. It surges to over 95% accuracy within the very first epoch and maintains this superior level with minimal variance. This rapid convergence and sustained stability ($\sim$10% gap over baselines) indicate that Co2S successfully establishes a high-quality supervision signal from the onset, thereby effectively stifling pseudo-label drift at its source and ensuring reliable optimization.
@@ IV-D Ablation Studies and Analysis
TABLE VII: Ablation study on the explicit and implicit semantic guidance mechanisms (ESG: Explicit Semantic Guidance; ISG: Implicit Semantic Guidance) on the WHDLD dataset.
@@ IV-D1 Impact of Explicit and Implicit Semantic Guidance
To validate the efficacy of the proposed explicit-implicit semantic co-guidance mechanism, we analyze the individual contributions of text embeddings and learnable queries which provide explicit and implicit class-level guidance respectively. As detailed in Table VII, the baseline model without guidance achieves 58.97 mIoU. Implementing implicit class-level guidance via learnable queries $\mathbf{Q}$ alone results in a slight performance drop to 58.86. This decline is attributed to the fact that the learnable queries are randomly initialized parameters lacking pre-trained priors. Optimizing these queries from scratch typically requires substantial annotated data, whereas the scarce labeled samples in our semi-supervised setting fail to provide sufficient supervision for their effective convergence. In contrast, integrating explicit class-level guidance via text embeddings $\mathbf{F}^{t}_{clip}$ alone yields a substantial gain of 1.8 and reaches 60.77 mIoU. This improvement stems from the inherent text-image alignment capability of CLIP acquired during pre-training, which enables the text embeddings to serve as powerful and reliable semantic anchors immediately. Notably, the optimal performance of 61.09 is attained only when both mechanisms are engaged. This result demonstrates a critical synergy where text embeddings ensure semantic correctness while learnable queries capture data-specific visual features, thereby jointly enhancing semantic consistency.
TABLE VIII: Ablation study on the heterogeneity of the dual-student architecture on the WHDLD dataset (1/24).
@@ IV-D2 Benefit from Global and Local Feature Fusion
To verify the effectiveness of the proposed global-local feature collaborative fusion strategy, we investigate the necessity of the heterogeneous dual-student design by comparing it against homogeneous initialization strategies, as detailed in Table VIII. First, the homogeneous configuration where both students utilize DINOv3 performs poorly, yielding an mIoU of only 45.20. This indicates that without strong global semantic priors, students initialized solely with local-structural features struggle to maintain category consistency, leading to severe semantic drift. Second, the homogeneous setting where both students employ CLIP attains a competitive mIoU of 60.78 due to rich global semantic representations. However, identical initialization results in highly correlated predictions. This lack of complementary viewpoints prevents the students from correcting each other’s mistakes, rendering them prone to coupled error accumulation and confirmation bias. In contrast, our proposed heterogeneous architecture combining CLIP and DINOv3 achieves the highest performance of 61.09 mIoU. This superiority confirms that the framework successfully synergizes the global semantic representations of CLIP with the local structural details of DINOv3. By arbitrating between these complementary feature streams, Co2S ensures both semantic correctness and boundary precision, thereby stabilizing the semi-supervised training under scarce data conditions.
TABLE IX: Ablation study on the contribution of different optimization objectives on the WHDLD dataset (1/24).
@@ IV-D3 Effectiveness of Optimization Objectives
To evaluate the individual contribution of each term in our overall training objective, we conduct an ablation study as listed in Table IX. The supervised baseline trained solely with supervised loss $\mathcal{L}_{\text{sup}}$ achieves 59.98 mIoU. Incorporating the consistency loss $\mathcal{L}_{\text{ct}}$ boosts the performance to 60.73, indicating that the weak-to-strong constraints effectively mine useful supervisory signals from unlabeled data to augment model training. Furthermore, adding the proposed stability loss $\mathcal{L}_{\text{sta}}$ yields the best performance of 61.09 mIoU. This further improvement confirms that $\mathcal{L}_{\text{sta}}$ is essential for mitigating pseudo-label drift by actively arbitrating between the two students based on confidence, thereby ensuring more reliable mutual learning.
TABLE X: Ablation study of synergizing CLIP with different SSM priors on the WHDLD dataset (1/24).
@@
     MAE [he2022masked]      60.84
     BEiTv2 [peng2022beit]      60.98
     iBOT [zhou2021ibot]      60.96
     SimMIM [xie2022simmim]      60.99
     DINOv3 [simeoni2025dinov3]      61.09
@@ IV-D4 Synergizing CLIP with Different SSM Priors
To assess the advantage of heterogeneous model collaboration, we evaluate CLIP [radford2021learning] paired with a range of self-supervised models as well as with itself as a homogeneous baseline. The baseline attains an mIoU of 60.78, representing the performance ceiling achievable by relying solely on vision-language priors. As reported in Table X, combining CLIP with any of the examined self-supervised models [he2022masked, peng2022beit, zhou2021ibot, xie2022simmim, simeoni2025dinov3] yields a performance gain over this baseline. The smallest improvement is 0.06 with MAE [he2022masked], while the largest improvement is 0.31 with DINOv3 [simeoni2025dinov3]. This consistent uplift across all pairings confirms that complementary visual priors from diverse self-supervised paradigms substantially enhance mutually semantic interactive learning beyond what CLIP alone can deliver.
@@ V Conclusion
[Uncaptioned image] Yi Zhou is pursuing the B.E. degree with the School of Computer Technology and Application at Qinghai University, Xining, China. His research interests primarily focus on computer vision, particularly in the field of remote sensing image processing.
[Uncaptioned image] Xuechao Zou received the B.E. degree in 2021 and the M.S. degree in 2024 from the School of Computer Technology and Application at Qinghai University, Xining, China. He is currently pursuing a Ph.D. degree with the School of Computer Science and Technology at Beijing Jiaotong University. His research interests focus on computer vision, particularly remote sensing image processing.
[Uncaptioned image] Shun Zhang is pursuing the B.E. degree with the School of Computer Technology and Application at Qinghai University, Xining, China. He is currently pursuing the M.S. degree with the School of Computer Science and Technology at Beijing Jiaotong University. His research interests primarily focus on computer vision, particularly in the field of remote sensing image processing.
[Uncaptioned image] Kai Li received his B.E. degree from the School of Computer Technology and Application at Qinghai University, Xining, China, in 2020, and his M.S. degree from the Department of Computer Science and Technology at Tsinghua University, Beijing, China, in 2024. He is pursuing a Ph.D. in the Department of Computer Science and Technology at Tsinghua University. His research interests focus primarily on speech separation and audio-visual learning.
[Uncaptioned image] Shiying Wang received the B.E. degree from Jilin Agricultural University, Changchun, China, in 2015, the M.S. degree and Ph.D. degree from the School of Computer Technology and Application, Qinghai University, Xining, China, in 2018 and 2025. Her research interests mainly include remote sensing image processing, particularly in grassland informatics.
[Uncaptioned image] Jingming Chen is pursuing the B.E. degree with the School of Computer Technology and Application at Qinghai University, Xining, China. His research interests primarily focus on computer vision, particularly in the field of remote sensing image processing.
[Uncaptioned image] Congyan Lang received the Ph.D. degree from the Beijing Key Laboratory of Traffic Data Analysis and Mining, School of Computer and Information Technology, Beijing Jiaotong University, Beijing, China, in 2006. She was a Visiting Professor with the Department of Electrical and Computer Engineering, National University of Singapore, Singapore, from 2010 to 2011. From 2014 to 2015, she was a visiting professor at the Department of Computer Science, University of Rochester, Rochester, NY, USA. She is a Professor at the School of Computer and Information Technology, Beijing Jiaotong University. She has published over 80 research articles in various journals and refereed conferences. Her research areas include computer vision and machine learning.
[Uncaptioned image] Tengfei Cao received the Ph.D. degree from Beijing University of Posts and Telecommunications (BUPT), Beijing, China, in July 2020. He is currently an Associate Professor with the Department of Computer Technology and Applications, Qinghai University, Xining, China. He has published several high-quality research papers, such as ACM Transactions on Multimedia Computing, Communications, and Applications, IEEE TRANSACTIONS ON NETWORK AND SERVICE MANAGEMENT, and CHINA COMMUNICATIONS. His research interests include multimedia communications and privacy security.
[Uncaptioned image] Pin Tao received his B.S. degree and Ph.D. in computer science and technology from Tsinghua University, Beijing, China, in 1997 and 2002. He is an Associate Professor at the Computer Science and Technology Department, Tsinghua University, Beijing, China. He is also the vice director of the Key Laboratory of Pervasive Computing, Ministry of Education. Dr. Tao has published more than 80 papers and over 10 patents. His current research interests mainly focus on human-AI hybrid intelligence and multimedia-embedded processing.
[Uncaptioned image] Yuanchun Shi received the PhD, MS and BS degrees in computer science from Tsinghua University. She is a Changjiang distinguished professor at the Department of Computer Science, Tsinghua University. She was a senior visiting scholar at the MIT AI Lab during 2001-2002. She had chaired several conferences, including ACM Ubicomp2011. She serves as the area editor of pervasive and mobile computing (PMC, Elsevier), editor of interacting with computer (IwC, Oxford University Press), and vice editor-in-chief of Communications of China Computer Federation (CCF). She has published more than one hundred papers in IJHCS, IEEE TPDS, TKDE, ACM CHI, MM, UIST, etc. Her research interests include human-computer interaction, pervasive computing, and multimedia communication. She is a Fellow of the IEEE.

### v2

@@ IV-C2 Qualitative Visualization Analysis
To further substantiate these quantitative findings, we provide a visual comparison of segmentation results across all six datasets in Fig. 4. As observed, baseline methods frequently struggle with semantic confusion in complex scenes. For instance, in the Potsdam dataset (Row f), UniMatch [47] and DWL [12] incorrectly classify large areas of impervious surface (white) as clutter (red). Similarly, in WHDLD (Row a), the vegetation (green) in the top-left corner is misidentified as water (blue) by most competitors. Moreover, our method demonstrates superior capability in handling fine-grained objects and texture details. In the MER dataset (Row c), baselines suffer from severe over-segmentation, generating excessive false positives for scattered small rocks (purple), whereas Co2S precisely suppresses this noise to match the Ground Truth. Likewise, in MSL (Row d), other methods tend to misclassify large areas of bedrock (blue) as gravel (olive green), while Co2S accurately preserves the correct material distribution. These visualizations confirm that our framework effectively harmonizes global semantic consistency with local structural precision across diverse and challenging remote sensing environments, successfully rectifying category confusion while preserving fine-grained details.
@@ IV-C3 Pseudo-label Quality and Stability
where $\hat{\mathbf{y}}_{u}$ is the pseudo-label, $\mathbf{y}_{gt}$ is the ground truth, and $c$ is the confidence score. The evolution curves for the first 10 epochs on the WHDLD dataset (1/24 split) are plotted in Fig. 5. The results reveal distinct training dynamics: 1) The supervised baseline (OnlySup) exhibits extreme volatility, oscillating drastically between 50% and 90%. This highlights the instability of learning from scratch with scarce annotations. 2) Standard methods like FixMatch [33] and UniMatch [33] show signs of early saturation. They quickly plateau at a sub-optimal accuracy level ($\sim$80%-88%), suggesting that they are trapped by confirmation bias early in the training process and fail to correct initial errors. 3) In contrast, Co2S (red curve) demonstrates immediate semantic alignment. It surges to over 95% accuracy within the very first epoch and maintains this superior level with minimal variance. This rapid convergence and sustained stability ($\sim$10% gap over baselines) indicate that Co2S successfully establishes a high-quality supervision signal from the onset, thereby effectively stifling pseudo-label drift at its source and ensuring reliable optimization.
@@ IV-D Ablation Studies and Analysis
TABLE VI: Ablation study on the explicit and implicit semantic guidance mechanisms (ESG: Explicit Semantic Guidance; ISG: Implicit Semantic Guidance) on the WHDLD dataset.
@@ IV-D1 Impact of Explicit and Implicit Semantic Guidance
To validate the efficacy of the proposed explicit-implicit semantic co-guidance mechanism, we analyze the individual contributions of text embeddings and learnable queries which provide explicit and implicit class-level guidance respectively. As detailed in Table VI, the baseline model without guidance achieves 58.97 mIoU. Implementing implicit class-level guidance via learnable queries $\mathbf{Q}$ alone results in a slight performance drop to 58.86. This decline is attributed to the fact that the learnable queries are randomly initialized parameters lacking pre-trained priors. Optimizing these queries from scratch typically requires substantial annotated data, whereas the scarce labeled samples in our semi-supervised setting fail to provide sufficient supervision for their effective convergence. In contrast, integrating explicit class-level guidance via text embeddings $\mathbf{F}^{t}_{clip}$ alone yields a substantial gain of 1.8 and reaches 60.77 mIoU. This improvement stems from the inherent text-image alignment capability of CLIP acquired during pre-training, which enables the text embeddings to serve as powerful and reliable semantic anchors immediately. Notably, the optimal performance of 61.09 is attained only when both mechanisms are engaged. This result demonstrates a critical synergy where text embeddings ensure semantic correctness while learnable queries capture data-specific visual features, thereby jointly enhancing semantic consistency.
TABLE VII: Ablation study on the heterogeneity of the dual-student architecture on the WHDLD dataset (1/24).
@@ IV-D2 Benefit from Global and Local Feature Fusion
To verify the effectiveness of the proposed global-local feature collaborative fusion strategy, we investigate the necessity of the heterogeneous dual-student design by comparing it against homogeneous initialization strategies, as detailed in Table VII. First, the homogeneous configuration where both students utilize DINOv3 performs poorly, yielding an mIoU of only 45.20. This indicates that without strong global semantic priors, students initialized solely with local-structural features struggle to maintain category consistency, leading to severe semantic drift. Second, the homogeneous setting where both students employ CLIP attains a competitive mIoU of 60.78 due to rich global semantic representations. However, identical initialization results in highly correlated predictions. This lack of complementary viewpoints prevents the students from correcting each other’s mistakes, rendering them prone to coupled error accumulation and confirmation bias. In contrast, our proposed heterogeneous architecture combining CLIP and DINOv3 achieves the highest performance of 61.09 mIoU. This superiority confirms that the framework successfully synergizes the global semantic representations of CLIP with the local structural details of DINOv3. By arbitrating between these complementary feature streams, Co2S ensures both semantic correctness and boundary precision, thereby stabilizing the semi-supervised training under scarce data conditions.
TABLE VIII: Ablation study on the contribution of different optimization objectives on the WHDLD dataset (1/24).
@@ IV-D3 Effectiveness of Optimization Objectives
To evaluate the individual contribution of each term in our overall training objective, we conduct an ablation study as listed in Table VIII. The supervised baseline trained solely with supervised loss $\mathcal{L}_{\text{sup}}$ achieves 59.98 mIoU. Incorporating the consistency loss $\mathcal{L}_{\text{ct}}$ boosts the performance to 60.73, indicating that the weak-to-strong constraints effectively mine useful supervisory signals from unlabeled data to augment model training. Furthermore, adding the proposed stability loss $\mathcal{L}_{\text{sta}}$ yields the best performance of 61.09 mIoU. This further improvement confirms that $\mathcal{L}_{\text{sta}}$ is essential for mitigating pseudo-label drift by actively arbitrating between the two students based on confidence, thereby ensuring more reliable mutual learning.
TABLE IX: Ablation study of synergizing CLIP with different SSM priors on the WHDLD dataset (1/24).
@@
     MAE [9]      60.84
     BEiTv2 [25]      60.98
     iBOT [54]      60.96
     SimMIM [46]      60.99
     DINOv3 [32]      61.09
@@ IV-D4 Synergizing CLIP with Different SSM Priors
To assess the advantage of heterogeneous model collaboration, we evaluate CLIP [27] paired with a range of self-supervised models as well as with itself as a homogeneous baseline. The baseline attains an mIoU of 60.78, representing the performance ceiling achievable by relying solely on vision-language priors. As reported in Table IX, combining CLIP with any of the examined self-supervised models [9, 25, 54, 46, 32] yields a performance gain over this baseline. The smallest improvement is 0.06 with MAE [9], while the largest improvement is 0.31 with DINOv3 [32]. This consistent uplift across all pairings confirms that complementary visual priors from diverse self-supervised paradigms substantially enhance mutually semantic interactive learning beyond what CLIP alone can deliver.